OpenStudy (anonymous):

A wheel made 517 revolutions at 34 4/9 revolutions per minute. How many minutes did it rotate?

OpenStudy (freckles):

\[\frac{a}{b} \frac{\text{ rev }}{\text{ min }} \\ \text{ I like to think this means } a \text{ rev } \text{ R }b \text{ min } \\ \text{ treating } R \text{ like a equal sign }\] now we want to know x such that we have 517 rev that we want to find x in 517 rev R x min ... \[a \text{ rev } \text{ R } b \text{ min} \\ \text{ divide both sides by } a \\ 1 rev R \frac{b}{a} \text{ min} \\ \text{ multiply both sides by } 517 \\ 517 \text{ rev }R \frac{517 \cdot b}{a} \text{ min }\] however this is the way I used to solve it you can also setup equivalent ratios and solve that way \[\frac{a}{b} =\frac{517}{x}\] and solve for x

OpenStudy (freckles):

I just used a/b for your 34 4/9 mixed fraction thing you may want to convert that to improper fraction

OpenStudy (anonymous):

\[\frac {rev} {time} = 34 + \frac {4}{9 } \] \[\frac {517} {time} = 34 + \frac {4}{9 } \]

OpenStudy (anonymous):

You guys are awesome, but I just can't get it, its not clicking in my head ):

OpenStudy (anonymous):

I keep getting an off number. i don't think i'm simplifying correctly. I keep getting 17807 7/9

OpenStudy (freckles):

\[34 \frac{4}{9} =\frac{34(9)+4}{9} =\frac{310}{9} \\ \text{ so you have } \\ \frac{310}{9}=\frac{517}{x}\]

OpenStudy (freckles):

can you solve that for x?

OpenStudy (anonymous):

4653/310?

OpenStudy (freckles):

yes

OpenStudy (freckles):

that says 310 rev per 9 min is equal to 517 rev per x min where we found x to be 4653/310 which is approximately 15 min

OpenStudy (anonymous):

I had that number! And when i tried to convert it to a mixed fraction it all got messed up! lol thank you so much! I'm currently 8 months pregnant and my child is eating up my brains. ):

OpenStudy (freckles):

oh you want to convert that to mix fraction?

OpenStudy (anonymous):

yes i got 15 3/310

OpenStudy (freckles):

alright sounds cool to me

OpenStudy (freckles):

I have no idea what it is like to be pregnant but I'm sure it isn't easy.

OpenStudy (anonymous):

No its not and even worse when you're taking online classes cause you can't get help on questions like this /: thanks again freckles!

OpenStudy (freckles):

np good luck and just one more month to go go you! you can do it! :)

OpenStudy (freckles):

unless you are an alien and you have to carry the baby for like 21 months or something then that is more sad

OpenStudy (anonymous):

Yes! I'm excited (: Oh no I'm not an alien, just human ! 9 months and thats it! I couldnt do it any longer haha

OpenStudy (freckles):

well thank goodness for that

OpenStudy (anonymous):

(:

OpenStudy (freckles):

anyways good luck with your math too (and congrats on the kid on the way)